OpenStudy (anonymous):

12y+8 = 5y+2 What is the answer every time I do it will not work?

OpenStudy (solomonzelman):

Subtract 5y from both sides and tell me what you get.

OpenStudy (anonymous):

7y+8=2

OpenStudy (solomonzelman):

yes, very nice!

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

Now, you would like to isolate the term that contains "7y", right? But, you have +8 on the same side. What should you do (to/from both sides) to get 7y by itself?

OpenStudy (anonymous):

subtract 8

OpenStudy (anonymous):

7y=-6

OpenStudy (solomonzelman):

Yes, subtract 8 from both sides, and this gives you?

OpenStudy (anonymous):

Am I not right @solomonzelman

OpenStudy (solomonzelman):

I am not trying to take over, but doing user's work is inappropriate according to the site's policy.

OpenStudy (anonymous):

Please am just trying to learn also

OpenStudy (anonymous):

its cool guys

OpenStudy (anonymous):

I can back off if u want me to

OpenStudy (solomonzelman):

yes, 7y=-6 is correct.

OpenStudy (anonymous):

its ok with me i was just confused for a little bit there on what you said daniel56k

OpenStudy (solomonzelman):

Now, you want to finish solving for y. y is multiplied times 7, (so you have 7 times y), and you want to get y on its own. What do you do to both sides to get rid of that 7 multiplied times y?

OpenStudy (anonymous):

divide

OpenStudy (solomonzelman):

Yes, and what would you divide both sides by, to get rid of the 7 multiplied on y?

OpenStudy (anonymous):

a really long decimal number but how is that a variable?

OpenStudy (solomonzelman):

we have 7y=-6 try to divide both sides by 7 (is tat what you did?)

OpenStudy (solomonzelman):

you can write the right side as a fraction

OpenStudy (anonymous):

yea y= 0.855714286

OpenStudy (solomonzelman):

but you had a NEGATIVE 6

OpenStudy (anonymous):

yea i still get same answer

OpenStudy (solomonzelman):

-6/7 = -0.855714286

OpenStudy (solomonzelman):

you get negative that, because you are dividing negative by positive.

OpenStudy (solomonzelman):

7y = -6 we divided both sides by 7, 7y ÷ 7 = -6 ÷ 7 and we get, y = -6/7

OpenStudy (solomonzelman):

I am writing the answer as a fraction, because that is more precise, more convenient. just better. You don't have to find the decimal equivalent to the fraction.

OpenStudy (anonymous):

ok what i am not getting is why there is a decimal as a variable

OpenStudy (solomonzelman):

Maybe you meant - why we get a solution for y, that is a decimal?

OpenStudy (solomonzelman):

(your wording is a little hard to understand)

OpenStudy (anonymous):

ok what i am not getting is why there is a decimal as a variable

OpenStudy (anonymous):

yea thats what i meant sorry

OpenStudy (solomonzelman):

There isn't anything wrong with that. You get a decimal/fraction.

OpenStudy (solomonzelman):

If you get a decimal/fraction, whether negative or positive, then that isn't an automatically faulty result; you know what I mean? y = some decimal is not an indication of an error

OpenStudy (solomonzelman):

And why we get a decimal, is because we have arrived at that solution after preforming the proper steps. Do you have any questions about the steps we did to solve the problem?

OpenStudy (anonymous):

no so the decimal is the answer and if so how would you simplify it?

OpenStudy (solomonzelman):

Example: \(\color{#000000 }{ \displaystyle 5z-9=3z+10 }\) Subtract \(3z\) from both sides, \(\color{#000000 }{ \displaystyle 5z-9\color{red}{-3z}=3z+10\color{red}{-3z} }\) I know that: 5z-3z=2z 3z-3z=0 (3z on the right cancel themselves out) And you get the following equation, \(\color{#000000 }{ \displaystyle 2z-9=10 }\) Now, I will add 9 to both sides, \(\color{#000000 }{ \displaystyle 2z-9\color{red}{+9}=10\color{red}{+9} }\) I know that: 10+9=19 -9+9=0 (9 on the left cancel themselves out) And you get, \(\color{#000000 }{ \displaystyle 2z=19 }\) then, we will divide both sides by 2, to solve for z. \(\color{#000000 }{ \displaystyle 2z\color{red}{\div 2}=19\color{red}{\div 2} }\) and I get, ((improper fraction)) \(\color{#000000 }{ \displaystyle z=19/2 }\) (it's called improper fraction becase numerator is bigger then denominator) you can also record your result as, ((decimal)) \(\color{#000000 }{ \displaystyle z=9.5 }\) and also, ((mixed fraction)) \(\color{#000000 }{ \displaystyle z=9\frac{1}{2} }\)

OpenStudy (solomonzelman):

@clarkgrace you know about fractions, correct? (look at the example later, please not now)

OpenStudy (anonymous):

to late

OpenStudy (solomonzelman):

what do you mean? you got to go?

OpenStudy (anonymous):

oh no i had already looked at the example problems